OpenStudy (anonymous):

sum1 help with my test asap Write the equation of a line in slope-intercept form that has a slope of -2 and passes through the point (2, -8). (Points : 2) y = -2x - 4 y = -2x - 8 y = -2x + 18 y = -2x – 12

OpenStudy (anonymous):

@Dido525

OpenStudy (anonymous):

Helping you on a test is something we cannot do here.

OpenStudy (anonymous):

its not relly test but its long like a test i jus thought if i sid test ppl wouldctuallly help me

OpenStudy (anonymous):

@Dido525

OpenStudy (anonymous):

@AriPotta heyycan u help plz

OpenStudy (anonymous):

@AccessDenied can u help please

OpenStudy (aripotta):

put the information into pint-slope form first

OpenStudy (aripotta):

point-slope*

OpenStudy (anonymous):

y=mx+b?

OpenStudy (aripotta):

that's slope-intercept form

OpenStudy (aripotta):

point-slope is y - y1 = m(x - x1)

OpenStudy (aripotta):

what do you get after putting the info in?

OpenStudy (anonymous):

so y-8=-2(x-2)?

OpenStudy (aripotta):

y + 8 = -2(x - 2)

OpenStudy (anonymous):

y+8=-2x+4?

OpenStudy (aripotta):

yes, and now "solve for y" to get it into slope-intercept form

OpenStudy (anonymous):

my bad -

OpenStudy (aripotta):

what did you do wrong?

OpenStudy (anonymous):

oh nvm so it would = y=-2x-4 bc u add 8

OpenStudy (aripotta):

yes

OpenStudy (anonymous):

ok wat about this one Rewrite the equation of a line in point-slope form to slope-intercept form:

OpenStudy (anonymous):

@AriPotta

OpenStudy (aripotta):

no need to tag me, i'm right here ;)

OpenStudy (anonymous):

:)

OpenStudy (aripotta):

right now it's in point-slope form. so do the same thing we did to the last one for this one

OpenStudy (anonymous):

y-3 wait its a fraction idk wat to do those confudse me

OpenStudy (aripotta):

for the last one, we distributed, correct? just do the same thing with this one. 1/2 times x is 1/2x and 1/2 times -8 is..?

OpenStudy (anonymous):

1/2-8?

OpenStudy (aripotta):

no. :/ what's half of 8?

OpenStudy (anonymous):

oh 4

OpenStudy (aripotta):

-4

OpenStudy (anonymous):

oh yea

OpenStudy (aripotta):

continue.

OpenStudy (anonymous):

=1/-2?

OpenStudy (aripotta):

ok. let's take this by steps starting with what we started with y - 3 = 1/2(x - 8) y - 3 = 1/2x - 4

OpenStudy (aripotta):

we need to continue to get into slope-intercept

OpenStudy (aripotta):

one step left

OpenStudy (anonymous):

+3 n +3 so y=1/2+1

OpenStudy (anonymous):

-1!

OpenStudy (aripotta):

yes, y = 1/2x - 1

OpenStudy (anonymous):

ok wat about this one now 2. Write the equation of a line in slope-intercept form passing through (4, 5) and (6, 8)

OpenStudy (aripotta):

make a new question thread

OpenStudy (aripotta):

please

OpenStudy (anonymous):

kk